OpenStudy (anonymous):

what is the distance between A(-8,4) and B(4,-1)?

jimthompson5910 (jim_thompson5910):

use the distance formula to get d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((4-(-8))^2+(-1-4)^2) d = sqrt((4+8)^2+(-1-4)^2) d = sqrt((12)^2+(-5)^2) d = sqrt(144+25) d = sqrt(169) d = 13 So exact distance between the two points is 13 units.